OpenStudy (anonymous):

Please take a quick look and see if i got the right answer. I just need help on this last question! What is the volume of a triangular prism with a height of 7 inches and a base that is an equilateral triangle with sides of 6 inches?

OpenStudy (anonymous):

Im not sure but i got 126in^3

OpenStudy (anonymous):

base area = \[\frac{1}{2} \sin 60 \times 6 \times 6\] volume = base area * height i got 15.58 * 7 =109.2 inch....ur wrng :(

OpenStudy (anonymous):

Oh well that;s not one of my answer though ;/. The answer are... 147 in3 84sqrt3 in3 126 in3 63sqrt3 in3

OpenStudy (anonymous):

well then thers sumthng wrng in the question

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

The formula for finding the area of a pyramid is V=1/3BH So first we have to do the thing where you find the total for the base but idk how to do it when it says a base that is an equilateral triangle with sides of 6 inches?

OpenStudy (anonymous):

After you find the total perimeter of the base then you input the numbers into the formula. I never saw anything about sin beign needed into this problem

OpenStudy (dumbcow):

haha is it a pyramid or a triangular prism ?

OpenStudy (anonymous):

triangle so i made a mistake by saying V=1/3bh instead of V=1/2bh

OpenStudy (dumbcow):

not quite Area of triangle = 1/2bh Volume = Area of triangle * height of prism

OpenStudy (anonymous):

yeah i meant it's a triangular prism

OpenStudy (dumbcow):

thushananth did it correctly, but just didn't leave it in terms of sqrt3

OpenStudy (anonymous):

Oh ok so how do we do that?

OpenStudy (dumbcow):

its 63sqrt3

OpenStudy (anonymous):

oh ok thanks but if it's no trouble can you show me how to work this out?

OpenStudy (dumbcow):

 if you know trig then sin(60) = h/6 --> h = 6*sqrt3/2 = 3*sqrt3 if you don't know trig then use pythagorean thm h = sqrt(6^2 - 3^2) = sqrt(36-9) = sqrt(27) = 3*sqrt3 Area triangle = 1/2 *6 * 3sqrt3 = 9sqrt3 Volume = 9sqrt3 * 7 = 63sqrt3